An electron has a speed of v = 0.939 c. What is its total energy (rest energy plus kinetic energy)? (in MeV)
Answer #1

Here ,

speed of particle , v = 0.939 c

energy of electron = m * c^2

energy of electron = m0*c^2/sqrt(1 -(v/c)^2 )

converting the units to MeV

energy of electron = 9.1 *10^-31 * (3 *10^8)^2/(sqrt(1 - 0.939^2) * 1.602 *10^-13)

energy of electron = 1.488 MeV

the energy of electron is 1.488 MeV